OpenStudy (anonymous):

The perimeter of a rectangle is 16x^2 +12x +18/3x in inches.The length measures 2x^2 +6x +9/3x in inches. Find an expression to represent the measure of the width in inches.

OpenStudy (mertsj):

Would you agree that perimeter of rectangle is 2L+2W?

OpenStudy (anonymous):

yes, p=2L+2w

OpenStudy (mertsj):

So replace P& L with the given values and solve the resulting equation for W

OpenStudy (mertsj):

\[16x^2+12x+\frac{18}{3x}=2(2x^2+6x+\frac{9}{3x})+2W\]

OpenStudy (mertsj):

Now distribute the 2 on the right side.

OpenStudy (anonymous):

4x^2+12x+18/3+2W?

OpenStudy (anonymous):

2l +2w=16x^2 +12x +18/3x

OpenStudy (anonymous):

2w=- 2( 2x^2 +6x +9/3x)+16x^2 +12x +18/3x= 12x^2

OpenStudy (anonymous):

w= 6x^2

OpenStudy (anonymous):

very confusing, but I think I got it. Thankyou so very much

OpenStudy (anonymous):

just thought there were more steps by having to factor it first

OpenStudy (anonymous):

my answer is right isn't it?

OpenStudy (anonymous):

the answers to choose are: x, 2x, 4x, or 6x I got 6x

OpenStudy (mertsj):

6x^2 is the correct answer.